OpenStudy (anonymous):

Eliana is reading a 931-page book. She reads 20 pages per day in the book. If x represents the number of days Eliana has spent reading the book so far, which of the following equations can be used to find the number of pages remaining in the book for Eliana to read? y = 20 - 931x y = 20 + 931x y = 931 - 20x y = 931 + 20x

OpenStudy (sleepyjess):

What do you think or what can you eliminate?

OpenStudy (anonymous):

i think its d

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

You almost have it. Lets try a number a days for x. How about she has been reading her book for 5 days already. 931+20(5)=pages left to read

OpenStudy (anonymous):

so wat letter

OpenStudy (sleepyjess):

With equation D 20*5=100 931+100=1031 She can't have more than 931 pages left to read so what do you think it is now?

OpenStudy (anonymous):

i think a but im not shure still @sleepyjess

OpenStudy (sleepyjess):

Try putting in 5 for x and see which one makes sense. y = 20 - 931(5) y = 20 + 931(5) y = 931 - 20(5) A y=20-931(5) y=20-4655 y=-4635 She cannot have a negative amount of pages left to read so A is out. B y=20+931(5) y=20+4655 y=4675 She cannot have more pages left to read than she started with so B is out. C y=931-20(5) y=931-100 y=831 This is the only one that makes sense because she now has less pages left to read than she started with.